OpenStudy (nuccioreggie):

If (62)p = 610, what is the value of p? 2 3 4 5

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

@Nnesha

OpenStudy (nuccioreggie):

@Nnesha

OpenStudy (nuccioreggie):

@Nnesha

Nnesha (nnesha):

what do you think ?

Nnesha (nnesha):

\[\huge\rm (x^m )^n = x^{ m \times n}\] exponent rule!

OpenStudy (nuccioreggie):

I thought it was B

OpenStudy (nuccioreggie):

B @Nnesha

Nnesha (nnesha):

let m =2 and n=3 so m times n = 10 ??

OpenStudy (nuccioreggie):

@Nnesha

Nnesha (nnesha):

yes ?

OpenStudy (nuccioreggie):

C @Nnesha

Nnesha (nnesha):

how ?

Nnesha (nnesha):

\[\huge\rm (x^m)^n = x^{m \times n}\] multiply outside exponent by inside exponent of variable

OpenStudy (nuccioreggie):

Ohhh it D @Nnesha

Nnesha (nnesha):

a b c or d but how ?? give me an explanation :-)

OpenStudy (nuccioreggie):

Ok but what do you think can u explain the answer to me please

Nnesha (nnesha):

i already did you should multiply both exponent \[\huge\rm (p^2)^m = p^{2 \times m } = p^{10}\] what is m?

Nnesha (nnesha):

what you should multiply with 2 to get 10 ?

OpenStudy (nuccioreggie):

5

Nnesha (nnesha):

yes that's right! hope that helpz

OpenStudy (nuccioreggie):

5

Nnesha (nnesha):

yes..